# G3Flow: Generative 3D Semantic Flow for Pose-aware and Generalizable Object Manipulation

### v1

@@ 2.3 Diffusion Models for Imitation Learning
Diffusion models [31, 7] are a powerful class of generative models that model the score of the distribution (the gradient of the energy) rather than the energy itself [30, 27]. The key idea behind diffusion models is their iterative transformation of a simple prior distribution into a target distribution through a sequential denoising process. In robotics, diffusion-based policies [1, 26, 38, 22, 8, 2, 19, 12, 13, 17] have demonstrated impressive performance in learning complex manipulation skills from demonstrations. Recent works have explored various directions: 3D Diffusion Policy [38] combines 3D scene representations with diffusion objectives, ChainedDiffuser [36] focuses on trajectory generation between keyposes, and 3D Diffuser Actor [10] tackles joint keyposes and trajectory prediction. However, these approaches primarily operate on geometric representations without explicit semantic understanding, limiting their precision in terminal-constrained manipulation and generalization across object variations. Our work G3Flow addresses this limitation by leveraging Foundation Models to maintain accurate and consistent semantic information during dynamic interactions, enabling more precise and generalizable manipulation capabilities.
@@ 4.4 Ablation Study
To explore the advantages of our complete, dynamic, object-level semantic flow representation, we conduct the ablation study with conventional scene-level feature clouds.
@@
| DP w/ G3Flow | ( 22.6%) | ( 41.2%) |
| Generation Freq. | |
| DP w/ Scene-Level Feature | 10.52 Hz |
| DP w/ G3Flow | 42.92 Hz ( 309.0%) |
We selected the Shoe Place (T) and Dual Shoes Place (T) tasks for comparison because they require adjustments to the orientation of the shoe throughout the entire trajectory, which rely more on long-term semantic understanding. Additionally, object occlusions are designed into the tasks, posing a greater challenge for semantic comprehension. The result is shown in Tab. 3 and Tab. 4.
Ablation on Quality of Semantic Field. As demonstrated in Tab. 3, our approach achieves a 22.6% and 41.2% increase in success rates for the Shoe Place and Dual Shoes Place tasks, respectively. The experimental results indicate that our method outperforms the direct acquisition of scene-level feature clouds. This is because the images we input into the VFM are centered around the object, as stated in Sec. 3, thus the returned features focus more on the object’s semantic information itself. Scene-level feature clouds are often interfered with by irrelevant semantic information from the scene. Moreover, directly obtained feature clouds cannot handle occlusions that occur during object interactions. As shown in Figure 10, our method can still obtain the complete object semantic flow even under occlusion, whereas traditional methods fail to do so, and there is a noticeable loss of the shoe’s semantic information itself.
Ablation on Field Generation Frequency. Robotic manipulation tasks have stringent requirements for real-time performance. We test the model inference speed on a single GPU machine with an unloaded NVIDIA GeForce RTX 4090. According to Table 4, our flow generation speed is over four times the speed of traditional field generation. This is because our entire process only requires a single call to the VFM, with subsequent flow maintenance achieved using Pose Tracking, which is much more lightweight than the VFM and has a faster inference speed [35, 32].
4.5 Visualizations

### v2

@@ 2.3 Diffusion Models for Imitation Learning
Diffusion models [32, 7] are a powerful class of generative models that model the score of the distribution (the gradient of the energy) rather than the energy itself [31, 28]. The key idea behind diffusion models is their iterative transformation of a simple prior distribution into a target distribution through a sequential denoising process. In robotics, diffusion-based policies [1, 27, 39, 23, 8, 2, 20, 12, 13, 18, 14] have demonstrated impressive performance in learning complex manipulation skills from demonstrations. Recent works have explored various directions: 3D Diffusion Policy [39] combines 3D scene representations with diffusion objectives, ChainedDiffuser [37] focuses on trajectory generation between keyposes, and 3D Diffuser Actor [10] tackles joint keyposes and trajectory prediction. However, these approaches primarily operate on geometric representations without explicit semantic understanding, limiting their precision in terminal-constrained manipulation and generalization across object variations. Our work G3Flow addresses this limitation by leveraging Foundation Models to maintain accurate and consistent semantic information during dynamic interactions, enabling more precise and generalizable manipulation capabilities.
@@ 4.4 Ablation Study
| Method | Field-Gen Freq. | Decision-making Freq. |
| DP w/ Scene-Level DINOv2 | 10.52 Hz (75.5%) | 9.52 Hz (72.0%) |
| DP w/ D3Fields (GenDP) | 8.14 Hz (81.0%) | 6.89 Hz (79.8%) |
| DP w/ G3Flow | 42.92 Hz | 34.04 Hz |
Ablation on Efficiency. Robotic manipulation tasks have stringent requirements for real-time performance. We test the model inference speed on a single GPU machine with an unloaded NVIDIA GeForce RTX 4090. We tested the inference speed of DP with scene-level DINOv2, DP with D3Fields (GenDP), and DP with G3Flow in Tab. 3. Our method significantly outperforms the baselines, achieving a decision frequency of 34.04 Hz, nearly 6 times faster than GenDP, meeting the requirements of most real-time robotic manipulation.
Our pipeline generates digital assets once per object, even across multiple operations. After the initial DINOv2 feature extraction, subsequent semantic flow estimation relies on a lightweight pose-tracking network [36, 33], enabling real-time performance.
Ablation on Quality of Semantic Field. To explore the advantages of our complete, dynamic, object-level semantic flow representation, we conduct the ablation study with conventional scene-level feature clouds and D3Fields.
We selected the Shoe Place (T) and Dual Shoes Place (T) tasks for comparison because they require adjustments to the orientation of the shoe throughout the entire trajectory, which rely more on long-term semantic understanding. Additionally, object occlusions are designed into the tasks, posing a greater challenge for semantic comprehension.
As shown in Table 4, our approach achieves a 22.6% and 41.2% increase in success rates for scene-level features in the Shoe Place (T) and Dual Shoes Place (T) tasks, respectively. For D3Fields features, the success rates improve by 9.3% and 3.7%, respectively. The experimental results demonstrate that D3Fields outperforms scene-level feature clouds, as it incorporates human prior knowledge, making the features more effective. However, both perform worse than G3Flow, which is due to the fact that the images we input into the VFM are centered around the object, as described in Sec. 3. This results in the returned features focusing more on the object’s semantic information. In contrast, scene-level feature clouds (including D3Fields, which are derived from scene-level point clouds through prior information processing) are often interfered with by irrelevant semantic information from the scene. Additionally, directly obtained feature clouds cannot handle occlusions that occur during object interactions. Our method, however, is able to obtain the complete object semantic flow even under occlusion, whereas scene-level features, including those from D3Fields, fail to do so, leading to a degradation in feature quality.
@@
| DP w/ D3Fields (GenDP) | ||
| DP w/ G3Flow |
In Fig. 9, we compare the feature quality of our method with that of scene-level feature and D3Fields. It can be observed that due to occlusions and background interference, scene-level features (D1) fail to distinguish the shoe’s toe and heel, which is crucial for adjusting the shoe’s pose. D3Fields benefits from goal image priors (C1), which improves semantic preservation over raw DINOv2. However, it has two drawbacks: (1) occluded regions (e.g., gripper coverage, C2) suffer from inaccurate semantics, and (2) it cannot recover a complete semantic point cloud from a single view, limiting precise manipulation (C3). In contrast, G3Flow generates a complete and high-quality semantic field by digital twin (B2).
4.5 Visualizations of G3Flow